OpenStudy (anonymous):

Please help!! Which of the following is a polynomial with roots 4, −5, and 7? f(x) = x3 − 6x2 − 27x + 140 f(x) = x3 − 6x2 − 20x + 27 f(x) = x3 − 20x2 − 27x + 35 f(x) = x3 − 20x2 − 35x + 140

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm~A~polynomial~with~root~}a~{\rm is}:}\) \(\large\color{black}{ \displaystyle f(x)=x-a}\) (A line) Same way, \(\large\color{black}{ \displaystyle {\rm~A~polynomial~with~roots~}a,~{\rm and}~b~{\rm is}:}\) \(\large\color{black}{ \displaystyle f(x)=(x-a)(x-b)}\) AND, same way \(\large\color{black}{ \displaystyle {\rm~A~polynomial~with~roots~}a,~b,~{\rm and}~c~{\rm is}:}\) \(\large\color{black}{ \displaystyle f(x)=(x-a)(x-b)(x-c) }\) And so on

OpenStudy (solomonzelman):

But, these are polynomials "of the least degree" and with "scale factor" of 1.

OpenStudy (solomonzelman):

It seems like your polynomials are indeed least degree polynomial. So just you my first post to find your polynomial.

OpenStudy (anonymous):

So how would I get x. Sorry I'm really confused:(

OpenStudy (solomonzelman):

??

OpenStudy (anonymous):

How would I start the problem? I read the lesson, but I don't get it

OpenStudy (solomonzelman):

See my first post? Look at the last polynomial with roots a, b, and c. The reason that this works, because my f(x)=0, when evaluated at x=a, b or c. ((Because 0 times anything is zero)) So you just model the same thing, but your a, b, and c are: 4, -5, and 7 (The order doesn't matter)

OpenStudy (solomonzelman):

`STEP 1:` Model the f(x) ((just as I showed, WITH a, b, and c BUT, YOU USE YOUR NUMBERS)) `Step 2` Then, EXPAND your polynomial.

OpenStudy (anonymous):

(x-4)(x-4)(x-7)? So then I would expand?

OpenStudy (solomonzelman):

check parenthesis #2, please.

OpenStudy (anonymous):

whoops. (x-4)(x-5)(x-7)?

OpenStudy (solomonzelman):

your second zero is NEGATIVE 5.

OpenStudy (solomonzelman):

(\(x-\color{red}{-5}\))=(x+5)

OpenStudy (anonymous):

(x-4)(x+5)(x-7)

OpenStudy (solomonzelman):

Yes, then expand :)

OpenStudy (anonymous):

So then I'd multiply?

OpenStudy (solomonzelman):

Yes, expand it...

OpenStudy (anonymous):

Would it be (x) = x^3 − 20x^2 − 35x + 140

OpenStudy (solomonzelman):

only the +140 and x^3 are correct... NO, it would not.

OpenStudy (solomonzelman):

next time, please show your work.

OpenStudy (solomonzelman):

Ok, what is (x-4)(x+5) ?

OpenStudy (solomonzelman):

(without showing work, I can't tell you where you made the error in expanding)

OpenStudy (anonymous):

x^3 − 6x^2 − 27x + 140

OpenStudy (solomonzelman):

Yes, perfect!

OpenStudy (anonymous):

I understand now:) Thanks!

OpenStudy (solomonzelman):

Alrighty ... yw